Flamo:

What is the radius of the circle whose equation is x^2 + y^2 = 8?

Flamo:

I'm Lost.

Angle:

so the equation is x^2 + y^2 = (radius)^2 so with the given equation (radius)^2 = 8 right?

Flamo:

Ye?

Angle:

so square root both sides to find the radius because square root of (radius)^2 = (radius)

Flamo:

Radius = 8?

Angle:

mmm square root

Flamo:

So 8^2?

Flamo:

x^2 + y^2 = 8^2?

Angle:

(radius)^2 = 8 so (radius) = square root of 8

Flamo:

So Radius is 4 because that's the square root of 8?

Angle:

are you sure the square root of 8 is 4?

Flamo:

Well, Isin't it \[2\sqrt{2}\]?

Angle:

that's right :)

Flamo:

So its x^2 + y^2 = \[2\]^2?

Flamo:

o, That didn't work

Angle:

hmmm I was gonna say that 2 sqrt(2) is the answer

Flamo:

O, Thats one of the answer Choices, I guess its Right

Flamo:

Ty!